the length of a triangle is one foot more than twice the width. If the perimeter is 242 ft. Find the dimensions.
Answer by checkley77(12844)

L=2W+1
2L=2W=242
2(2W+1)+2W=242
4W+2=2W=242
6W=242-2
6W=240
W=240/6
W=40 FT. ANS.FOR THE WIDTH.
L=2*40+1
L=80+1=81 FT. ANS. FOR THE LENGTH.
PROOF:
2*81+2*40=242
162+80=242
242=242